OpenStudy (jojo12g):

Use a transformation to solve the equation. f – 12 = –35 A. f = –47 B. f = –23 C. f = 23 D. f = 47

OpenStudy (agl202):

F - 12 + 12 = -35 +12 or f =-23 or B